# Quantum dot energy levels in bilayer graphene: Exact and approximate study

### v1

@@ I Introduction
The versatile electronic properties of bilayer graphene (BLG) [1, 2], along with experimental advancements have enabled the fabrication of highly controllable quantum devices which might find applications in emerging quantum technologies. Quantum dots (QDs) are among these quantum devices and various studies have explored control and manipulation of confined charge, valley, and spin states [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In BLG confined quantum states can be defined by electrostatic gates providing a simple and attractive platform to realize tunable quantum dots [14, 15]. A perpendicular magnetic field can lift the valley degeneracy offering additional control of the discrete energy levels and quantum states of the QD. The resulting QD can be used to host qubits [8, 9] which can be efficiently controlled with external electric and magnetic fields. Beyond the spectroscopic investigation and characterization of single QDs in BLG [12], coherent charge oscillations have been successfully demonstrated [16] and measurements of valley [7] and spin lifetimes [3] have been performed.
@@
We describe in Sec. II the physical model of the QD in BLG based on the four-component continuum model. We continue in Sec. III to detail the necessary steps that allow us to derive the approximate QD equations. In Sec. IV we identify with specific examples some QD regimes that can be studied approximately. In Sec. V a detailed comparison is made between the exact and the approximate QD energies demonstrating the accuracy of our method. The conclusions of the work are summarized in Sec. VI, and some further results are presented in Appendix A.
@@ III Derivation of approximate quantum dot equations
Equations (10a) and (10b) are exact, and to proceed with some approximations we set
@@
and assume the function $\lambda$ to be approximately constant in the region where the QD states are confined. This regime has been shown to be valid for the Landau states [19] but we expect deviations from this regime when $V_{W}\neq 0$, since the resulting confining region of the QD states is now modified. However, these deviations are small when the character (profile) of the Landau states is not significantly altered due to $V_{W}$ [18], therefore $\lambda$ can be assumed to be approximately constant beyond the Landau regime. This case has been demonstrated for different potential wells defining QDs in monolayer graphene [18, 20]; consequently, a similar framework should be valid in bilayer graphene. A characteristic example of small deviations from the Landau regime due to $V_{W}\neq 0$ involves the creation of an extra node in the induced states [21]. This node lies in a region where the amplitude of the states is exponentially small, and as a result the creation of the node only weakly affects the character of the Landau states [21].
@@
Equation (20) can be slightly simplified in the limit where $q_{1}\approx q_{2}$. The validity of this limit can be demonstrated by starting with Eqs. (10a) and (10b) and performing the transformation
@@
which leads to the following coupled equations
@@
Equations (22a) and (22b) are exact and suggest that $q_{2}\approx q_{1}$ when $\chi_{+}=w+\nu\approx 0$, so that both sides of Eq. (22a) simultaneously vanish. As a result, with $\chi_{-}=2\nu$ the derived approximate equation for $\nu$ becomes a second order equation identical to Eq. (19), but with $Q_{+}$ replaced by $\tilde{Q}_{+}$. Our calculations show that $Q_{+}$ gives more accurate approximate results and for this reason we use $Q_{+}$ throughout this work.
@@
while “excited” states can be written in terms of Hermite polynomials [22]. The number of nodes of a QD state in monolayer graphene is sensitive to the position of the QD level relative to the potential well [21]. It is straightforward to show that this relativistic effect occurs also in a bilayer graphene QD but is not captured by the approximate Eq. (25). Consequently, within our approximate model a QD state, when $V_{\rm QD}\neq 0$, has the same number of nodes as the original Landau state. We further address this point in Sec. V.
@@ IV Approximate quantum dot regimes
Figure 1(c) presents $Q_{+}$ for a typical case where anticrossing points can be observed in the QD energy spectrum for energies $E<V_{\rm asym}$ (Appendix A). Assuming $V_{\rm b}=0$ the function $Q_{+}$ has now two positive maxima separated by a central region where the singular point at $r\approx 44$ nm dominates. According to the above remarks the approximate method developed in Sec. III becomes less accurate as $V_{W}$ increases and the QD levels start to form anticrossings. The transition from the regime shown in Fig. 1(a) to that in Fig. 1(c) can be observed, for example, by increasing $V_{W}$. Between the two regimes Eq. (25) may still be used, provided the inner maximum of $Q_{+}$ dominates. This implies that the QD state has negligible amplitude around the outer maximum, hence, this configuration is more likely to occur for small values of $V_{W}$ (Sec. V).
@@
Figure 3: Exact and approximate QD energy levels as a function of QD potential well for $L_{W}=100$ nm, $B=4$ T, $s=2$, and $V_{\rm b}=0$. From lower to upper curve: $m=-4$, $-16$, $-25$, $-35$, $-45$. Approximate QD levels are derived from Eq. (25) for different $n$ values.
@@ V Comparison between exact and approximate energies
For the numerical calculations of the exact QD levels we employ a similar numerical approach to that in Ref. 23. Some simple modifications are needed to apply the appropriate boundary conditions that result in a Hermitian eigenvalue problem [23]. In the present work the derivatives in the radial Eqs. (5a)–(5d) are approximated by a second-order finite-difference scheme with a discretization length $\delta$. Compared with the first-order scheme employed in Ref. 23 the advantage is that the numerical errors are of the order of $\delta^{2}$. The disadvantage is that the resulting matrix is less sparse, however, it can be easily treated using ARPACK subroutines incorporated into a python program. For the numerical calculations we take $\delta<0.2$ nm which ensures that the induced numerical errors are small and do not affect the comparison between the exact and the approximate QD levels.
@@
In Fig. 2 we only plot QD levels which emerge from positive bulk Landau levels, however, the approximate method can also be successfully applied to QD levels that originate from negative Landau levels (Appendix A). Similar to monolayer graphene [20], by increasing $V_{W}$ some QD energy levels start to anticross for energies $E<V_{\rm asym}$. The resulting QD levels arise from the hybridization between electron and hole Landau levels, and in this regime the function $Q_{+}$ has the characteristic form shown in Fig. 1(c). As discussed in Sec. III this regime is beyond the scope of the present work and cannot be captured by Eq. (25).
Figure 3 presents some QD energy levels as a function of the QD potential well for negative $m$ values. The agreement between the approximate energy levels and the exact ones is again very good especially for small values of $V_{W}$. Quantum dot levels can also emerge from the zero-energy Landau level, however, for these levels the assumption, $1/\lambda=\nu/w$, [Eq. (12)] is not satisfied and Eq. (25) cannot be applied. In this QD regime the singular point is expected to be important and a different strategy is needed to obtain the approximate QD energies. However, we can still obtain some physical insight without significantly deviating from the analysis in Sec. III. We can loosely assume that $\phi_{A}\gg\phi_{B^{\prime}}$ and then set $\chi_{\pm}\approx w$ in Eqs. (22a) and (22b) to derive a single differential equation for $w$. The resulting equation is similar to Eq. (19) with $Q_{\pm}(r)=q_{1}(r)\pm[E-V_{\rm QD}(r)]t_{c}/\gamma^{2}$ when $V_{\rm b}=0$. This approximate equation is particularly good for small values of $m$ and $V_{W}$. As an example, at $V_{W}=31$ meV the relative error is about $0.5\%$ for $m=-4$ but this quickly increases to $5\%$ for $m=-16$.
The results in Fig. 2 and Fig. 3 are for $s=2$ and we have confirmed that good agreement between exact and approximate energies is found for different values of $s$ (Appendix A). We briefly explore the role of $s$ in Fig. 4, where we plot the QD levels as a function of the potential well width $L_{W}$. Here, the potential well is fixed and equal to $V_{W}\approx E_{\rm L}$, so for this reason the QD energy tends to $E\approx 0$ with $L_{W}$. This behaviour is robust and agrees with the approximate Eq. (27). This confirms the fact that the decrease of a QD level from the bulk Landau level cannot be greater than $V_{W}$, and according to Fig. 4 for larger $s$ values the QD energy saturates at smaller potential widths.
@@
The accuracy of the harmonic oscillator formula, Eq. (25), depends on how well the parabolic term can approximate $Q_{+}$ in Eq. (24). This, however, cannot be predicted in advance (Appendix A) due to the many terms involved in the approximate equations, as well as the broad range of the various physical parameters. Our numerical investigation suggests that the approximate results are usually less accurate for negative angular momenta and, in general, for smaller absolute values of angular momenta. For small energies, especially near zero, the relative errors can be large but the correct characteristics can still be predicted. In contrast, the relative errors tend to be smaller at higher magnetic fields and larger QD energies as well as for larger values of $L_{W}$. Additionally, the accuracy of the approximate QD levels derived from Eq. (25) can be improved by numerically solving the approximate Eq. (19).
VI Conclusion
@@
We proved the efficacy of the approximate method by comparing the approximate energies with the exact energies derived from the four-component continuum model in BLG. We found that the approximate energy levels as a function of the magnetic field and QD potential well exhibit the correct general features, such as for example, anticrossing points as well as a linear dependence on the dot potential, and agree very well with the exact energy levels in a broad range of QD parameters. We demonstrated various realistic regimes where the relative error can be vanishingly small. Semiclassical and quantum super-symmetric methods might be interesting alternative frameworks to approximately explore QD physics in bilayer graphene.
@@
Appendix A Additional examples of quantum dot energies

### v2

@@ I Introduction
The versatile electronic properties of bilayer graphene (BLG) [1, 2], along with experimental advancements have enabled the fabrication of highly controllable quantum devices which might find applications in emerging quantum technologies. Quantum dots (QDs) are among these quantum devices and various studies have explored control and manipulation of confined charge, valley, and spin states [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In BLG confined quantum states can be defined by electrostatic gates providing a simple and attractive platform to realize tunable quantum dots [12, 15]. A perpendicular magnetic field can lift the valley degeneracy offering additional control of the discrete energy levels and quantum states of the QD. The resulting QD can be used to host qubits [15, 7] which can be efficiently controlled with external electric and magnetic fields. Beyond the spectroscopic investigation and characterization of single QDs in BLG [10], coherent charge oscillations have been successfully demonstrated [16] and measurements of valley [6] and spin lifetimes [3] have been performed.
@@
We describe in Sec. II the physical model of the QD in BLG based on the four-component continuum model. We continue in Sec. III to detail the necessary steps that allow us to derive the approximate QD equations. In Sec. IV we identify with specific examples some QD regimes that can be studied approximately. In Sec. V a detailed comparison is made between the exact and the approximate QD energies demonstrating the accuracy of our method. The conclusions of the work are summarized in Sec. VI and some further results are presented in two Appendixes.
@@ III Derivation of approximate quantum dot equations
Equations (10a) and (10b) are exact and to proceed with some approximations we set
@@
and assume the function $\lambda$ to be approximately constant in the region where the QD states are confined. This regime has been shown to be valid for the Landau states [19] but we expect deviations from this regime when $V_{W}\neq 0$, since the resulting confining region of the QD states is now modified. However, these deviations are small when the character (profile) of the Landau states is not significantly altered due to $V_{W}$ [18], therefore $\lambda$ can be assumed to be approximately constant beyond the Landau regime. This case has been demonstrated for different potential wells defining QDs in monolayer graphene [18, 20, 21]; consequently, a similar framework should be valid in bilayer graphene. A characteristic example of small deviations from the Landau regime due to $V_{W}\neq 0$ involves the creation of an extra node in the induced QD states [22]. Provided this node lies in a region where the amplitude of the states is exponentially small it can only weakly affect the original character of the Landau states [22]. In Appendix A we further examine the condition $w=\lambda\nu$.
@@
For completeness we now follow a slightly different strategy to derive an approximate QD equation similar to Eq. (19). Starting with Eqs. (10a) and (10b) and performing the transformation
@@
leads to the following coupled equations
@@
Equations (22a) and (22b) are exact and suggest that $\chi_{+}=w+\nu\approx 0$ when $q_{2}\approx q_{1}$ so that both sides of Eq. (22a) simultaneously vanish. As a result, with $\chi_{-}\approx 2\nu$ the final approximate equation for $\nu$ becomes a second order equation identical to Eq. (19), but with $Q_{+}$ replaced by $\tilde{Q}_{+}$. Our numerical calculations demonstrate that in general $Q_{+}$ gives more accurate approximate results and for this reason we use $Q_{+}$ throughout this work. In Appendix A we further investigate the approximate model and clarify in a rigorous way how the condition $q_{2}\approx q_{1}$ is essentially related to the assumption that $\lambda$ is constant.
@@
while “excited” states can be written in terms of Hermite polynomials [23]. The number of nodes of a QD state in monolayer graphene is sensitive to the position of the QD level relative to the potential well [22]. It is straightforward to show that this relativistic effect occurs also in a bilayer graphene QD but is not captured by the approximate Eq. (25). Consequently, within our approximate model a QD state, when $V_{\rm QD}\neq 0$, has the same number of nodes as the original Landau state. We return to this point in Sec. V.
@@ IV Approximate quantum dot regimes
Figure 1(c) presents $Q_{+}$ for a typical case where anticrossing points can be observed in the QD energy spectrum for energies $E<V_{\rm asym}$ (Appendix B). Assuming $V_{\rm b}=0$ the function $Q_{+}$ has now two positive maxima separated by a central region where the singular point at $r\approx 44$ nm dominates. According to the above remarks the approximate method developed in Sec. III becomes less accurate as $V_{W}$ increases and the QD levels start to form anticrossings. The transition from the regime shown in Fig. 1(a) to that in Fig. 1(c) can be observed, for example, by increasing $V_{W}$. Between the two regimes Eq. (25) may still be used, provided the inner maximum of $Q_{+}$ dominates. This implies that the QD state has negligible amplitude around the outer maximum, hence, this configuration is more likely to occur for small values of $V_{W}$ (Sec. V).
@@
Figure 3: Exact and approximate QD energy levels as a function of QD potential well for $L_{W}=100$ nm, $B=4$ T, $s=2$, and $V_{\rm b}=0$. From lower to upper curve: m=4m=-4, 16-16, 25-25, 35-35, 45-45. Approximate QD levels are derived from Eq. (25) for different $n$ values.
@@ V Comparison between exact and approximate energies
For the numerical calculations of the exact QD levels we employ a similar numerical approach to that in Ref. 24. Some simple modifications are needed to apply the appropriate boundary conditions that result in a Hermitian eigenvalue problem [24]. In the present work the derivatives in the radial Eqs. (5a)–(5d) are approximated by a second-order finite-difference scheme with a discretization length $\delta$. Compared with the first-order scheme employed in Ref. 24 the advantage is that the numerical errors are of the order of $\delta^{2}$. The disadvantage is that the resulting matrix is less sparse, however, it can be easily treated using ARPACK subroutines incorporated into a python program. For the numerical calculations we take $\delta<0.2$ nm which ensures that the induced numerical errors are small and do not affect the comparison between the exact and the approximate QD levels.
@@
In Fig. 2 we only plot QD levels which emerge from positive bulk Landau levels, however, the approximate method can also be successfully applied to QD levels that originate from negative Landau levels (Appendix B). Similar to monolayer graphene [20], by increasing $V_{W}$ some QD energy levels start to anticross for energies $E<V_{\rm asym}$. The resulting QD levels arise from the hybridization between electron and hole Landau levels, and in this regime the function $Q_{+}$ has the characteristic form shown in Fig. 1(c). As discussed in Sec. III this regime is beyond the scope of the present work and cannot be captured by Eq. (25).
Figure 3 presents some QD energy levels as a function of the QD potential well for negative $m$ values. The agreement between the approximate energy levels and the exact ones is again very good especially for small values of $V_{W}$. Quantum dot levels can also emerge from the zero-energy Landau level, however, for these levels the assumption, $1/\lambda=\nu/w$, [Eq. (12)] is not satisfied and Eq. (25) cannot be applied. In this QD regime the singular point is expected to be important and a different strategy is needed to obtain the approximate QD energies. However, we can still obtain some physical insight without significantly deviating from the analysis in Sec. III. We can loosely assume that $\phi_{A}\gg\phi_{B^{\prime}}$ and then set $\chi_{\pm}\approx w$ in Eqs. (22a) and (22b) to derive a single differential equation for $w$. The resulting equation is similar to Eq. (19) with $Q_{\pm}(r)=q_{1}(r)\pm[E-V_{\rm QD}(r)]t_{c}/\gamma^{2}$ when $V_{\rm b}=0$. This approximate equation is particularly good for small values of $m$ and $V_{W}$. As an example, at $V_{W}=31$ meV the relative error is about $0.5\%$ for m=4m=-4 but this quickly increases to $5\%$ for m=16m=-16.
The results in Fig. 2 and Fig. 3 are for $s=2$ and we have confirmed that good agreement between exact and approximate energies is found for different values of $s$ (Appendix B). We briefly explore the role of $s$ in Fig. 4, where we plot the QD levels as a function of the potential well width $L_{W}$. Here, the potential well is fixed and equal to $V_{W}\approx E_{\rm L}$, so for this reason the QD energy tends to $E\approx 0$ with $L_{W}$. This behaviour is robust and agrees with the approximate Eq. (27). This confirms the fact that the decrease of a QD level from the bulk Landau level cannot be greater than $V_{W}$, and according to Fig. 4 for larger $s$ values the QD energy saturates at smaller potential widths.
@@
The accuracy of the harmonic oscillator formula, Eq. (25), depends on how well the parabolic term can approximate $Q_{+}$ in Eq. (24). This, however, cannot be predicted in advance (Appendix B) due to the many terms involved in the approximate equations, as well as the broad range of the various physical parameters. Our numerical investigation suggests that the approximate results are usually less accurate for negative angular momenta and, in general, for smaller absolute values of angular momenta. For small energies, especially near zero, the relative errors can be large but the correct characteristics can still be predicted. In contrast, the relative errors tend to be smaller at higher magnetic fields and larger QD energies as well as for larger values of $L_{W}$. Additionally, the accuracy of the approximate QD levels derived from Eq. (25) can be improved by numerically solving the approximate Eq. (19).
VI Conclusion and discussion
@@
We proved the efficacy of the approximate method by comparing the approximate energies with the exact energies derived from the four-component continuum model in BLG. We found that the approximate energy levels as a function of the magnetic field and QD potential well exhibit the correct general features, such as for example, anticrossing points as well as a linear dependence on the dot potential, and agree very well with the exact energy levels in a broad range of QD parameters. We demonstrated various realistic regimes where the relative error can be vanishingly small.
Semiclassical and quantum super-symmetric methods might be interesting alternative frameworks [25] to approximately explore QD physics in bilayer graphene. Equation (19) can be used as the starting point of a WKB method [23, 25] with $\sqrt{Q_{+}}$ defining the local wavelength. The approximate QD energies can be obtained from a standard Bohr-Sommerfeld quantization condition involving the classical turning points, $Q_{+}=0$. Each method is expected to have some advantages in terms of numerical effort, accuracy, and pedagogical insight. The harmonic oscillator formula derived in the present work is easily implementable and accurate enough. An interesting case which seems to remain unexplored corresponds to the quantum dot regime where a singular point separates two classically allowed regions [Fig. 1(c)]. This configuration that couples electron and hole levels cannot be captured by our approximate model. A more powerful QD model should treat equally well the different regimes illustrated in Fig. 1.
Acknowledgements.
I wish to thank the anonymous referees for some useful comments.
@@ Acknowledgements.
Appendix A Examining the approximate quantum dot model
In the derivation of the approximate QD model we consider that the QD states satisfy $w=\lambda\nu$, and examine the simplest (adiabatic) regime where the function $\lambda$ is to a good approximation constant. To further explore this aspect we start with Eqs. (13) and (14) and eliminate $\nu$ by using Eq. (19). We derive that
$\frac{1}{\lambda}=\frac{\gamma^{2}}{t_{c}\sqrt{\kappa}}(q_{1}-Q_{\pm}),$ (28)
and
$\lambda=\frac{\gamma^{2}}{t_{c}\sqrt{\kappa}}(q_{2}-Q_{\pm}),$ (29)
These two equations lead to
$\frac{t_{c}\sqrt{\kappa}}{\gamma^{2}}\lambda^{2}-(q_{2}-q_{1})\lambda=\frac{t_{c}\sqrt{\kappa}}{\gamma^{2}}.$ (30)
If $q_{1}\approx q_{2}$ then to a good approximation $\lambda$ is constant, $|\lambda|\approx 1$. This limit is particularly good for large $m$ and when the potential terms $V^{{}^{\prime}}_{j}$, $V^{{}^{\prime\prime}}_{j}$ are small. This means that the QD states should be confined inside a region where the QD potential varies slowly. Within a more general analysis, defining the $r$-dependent function
$\lambda_{0}(r)=\gamma^{2}\frac{q_{2}-q_{1}}{2t_{c}\sqrt{\kappa}},$ (31)
and taking the limit $|\lambda_{0}|\ll 1$ we obtain from Eq. (30)
$\lambda\approx\pm 1+\lambda_{0}(r)\pm\frac{\lambda^{2}_{0}(r)}{2}.$ (32)
Here, we assume that any singular points can be safely ignored and that the QD states are nicely confined within the single maximum or two maxima of $Q_{+}$ as shown in Fig. 1. Thus the classical turning points, $Q_{+}=0$, define the boundaries of the regions in which the QD states are confined. Provided $|\lambda_{0}|\ll 1$ we can treat $\lambda$ as constant and the approximate QD model is expected to be accurate.
In the approximate QD model, assuming that $w^{{}^{\prime\prime}}\approx\lambda\nu^{{}^{\prime\prime}}$ implies that the quantity
$\varepsilon=\lambda^{{}^{\prime\prime}}\nu+2\lambda^{{}^{\prime}}\nu^{{}^{\prime}}$ (33)
is vanishingly small ($\lambda$ is nearly constant) and is therefore discarded. This process induces an error. One possible way to estimate it is to use the state $\nu$ and energy $E$, initially computed for $\lambda=1$, to define the “correction” term $\delta q_{1}=\varepsilon/\lambda\nu$. This correction comes from the exact QD equations while $\lambda$ can be computed directly from Eq. (30). We can then replace in the approximate Eq. (13) $q_{1}$ with $q_{1}+\delta q_{1}$ and follow the same steps as in the main text. The final equation has the same form as Eq. (19) and has to be solved (exactly or perturbatively) again for $\nu$ and $E$.
In our approach an additional error is induced due to the expansion of $Q_{+}$ about the position of the maximum [Eq. (24)], and the reduction of the differential equation for $\nu$ to that of a harmonic oscillator. Therefore, to account simultaneously for both errors we directly compare exact and approximate energies for a few characteristic cases in Appendix B. We finally remark that the oscillator can also be used to drastically simplify $\delta q_{1}$ (at least) for the nodeless state $\nu$; since $\nu^{{}^{\prime}}/\nu=(r_{0}-r)Q_{+}(r_{0})$ only terms involving $\lambda$ are relevant.
Appendix B Additional examples of quantum dot energies